### v1

# Strong anharmonicity dictates ultralow thermal conductivities of type-I clathrates

Abstract
Type-I clathrate solids have attracted significant interest due to their ultralow thermal conductivities and subsequent promise for thermoelectric applications, yet the mechanisms underlying these properties are not well understood. Here, we extend the framework of vibrational dynamical mean-field theory (VDMFT) to calculate temperature-dependent thermal transport properties of $X_{8}$a16Ge30, where $X=$ Ba, Sr, using a many-body Green’s function approach. We find that nonresonant scattering between cage acoustic modes and rattling modes leads to a reduction of acoustic phonon lifetimes and thus thermal conductivities. Moreover, we find that the moderate temperature dependence of conductivities above 300 K, which is consistent with experimental measurements, cannot be reproduced by standard perturbation theory calculations, which predict a $T^{-1}$ dependence. Therefore, we conclude that nonperturbative anharmonic effects, including four- and higher-phonon scattering processes, are responsible for the ultralow thermal conductivities of type-I clathrates.
I Introduction
Good thermoelectric materials have high electronic conductivities but low thermal conductivities, which presents a challenging problem in materials design [1]. One promising class of materials are type-I clathrate solids, which contain cage-like chemical structures of covalently bonded atoms that host loosely bound “guest” atoms (Fig. 1a), although the mechanism responsible for their low thermal conductivity has been debated [2, 3, 4, 5, 6, 7]. The rattling displacements of the guest atoms generate dispersionless, low-frequency optical phonon branches that intersect the acoustic phonon branches of the cage lattice [3]. The hybridization of these branches suppresses the group velocity of the acoustic phonons, which also inherit the reduced lifetimes of the strongly anharmonic optical rattling phonons [4, 8]. More exotic behaviors have also been proposed, including a phonon Kondo effect [6].
A standard approach to calculate the thermal conductivity tensor $\kappa^{ij}$ is based on the Boltzmann transport equation in the single-mode approximation (BTE-SMA) [9],
$\kappa^{ij}_{\text{SMA}}=\frac{1}{V}\sum_{\bm{k}}\sum_{\lambda}C_{\lambda}(\bm{k})v^{i}_{\lambda}(\bm{k})v^{j}_{\lambda}(\bm{k})\tau_{\lambda}(\bm{k})\,,$ (1)
where $V$ is the cell volume, $\lambda$ is an index over phonon modes in the system, $C_{\lambda}(\bm{k})$ is the heat capacity of mode $\lambda$, and $v^{i}_{\lambda}(\bm{k})$ is the phonon group velocity of mode $\lambda$ in Cartesian direction $i$. The mode lifetimes $\tau_{\lambda}(\bm{k})$ are commonly obtained by treating three-phonon processes to second order in perturbation theory (PT). Given the strong anharmonicity required to suppress the lifetimes of acoustic phonons in materials with low thermal conductivities, it is natural to question the quantitative accuracy of this approach. In particular, at temperatures above the Debye temperature, the heat capacity is constant, and PT predicts that the lifetimes decrease as $\tau\propto T^{-1}$, such that $\kappa\propto T^{-1}$. Experimentally observed deviations from this temperature dependence are indicative of stronger anharmonicity [10, 11], and their theoretical explanation requires more advanced methods. Approaches relying on equilibrium and nonequilibrium molecular dynamics (MD) simulations [2, 12, 13, 14] include nonperturbative effects of strong anharmonic interactions, but the computational expense associated with such direct simulation, especially when considering finite-size effects, makes them intractable for systems with complex unit cells. Furthermore, obtaining systematic or mechanistic insight from such large-scale atomistic simulations can be challenging relative to approaches, such as the BTE-SMA, which calculate phonon-mode resolved thermal conductivities.
Recently, we introduced vibrational dynamical mean-field theory (VDMFT) as an affordable but nonperturbative computational method for simulating anharmonic lattice dynamics [15], and we used it to calculate the phonon spectral function of the type-I clathrates Ba8Ga16Ge30 (BaGG) and Sr8Ga16Ge30 (SrGG) [8]. Here, we extend the VDMFT methodology to enable the calculation of conductivities (neglecting vertex corrections) [16], which we then use to calculate the temperature-dependent conductivities of the same clathrate systems studied in our previous work. We also briefly evaluate the evidence for a phonon Kondo effect [6], which our method is uniquely suited to study, given the success of DMFT and its associated impurity problem in explaining Kondo physics in electronic systems.
Through our VDMFT calculations, we confirm that anharmonic interactions between acoustic modes of the cage lattice and flat optical modes corresponding to rattling motions of guest atoms inside 24-atom cages are responsible for the ultralow thermal conductivities of these materials. Furthermore, through a nonperturbative description of this anharmonicity, we find that the thermal conductivity has a relatively moderate dependence on the temperature above 300 K, in agreement with experimental observations. These results cannot be captured, even qualitatively, using conventional lowest-order PT, suggesting that nonperturbative effects of anharmonicity, including four-phonon and higher-order scattering processes, are key to understanding the reduced thermal transport of type-I clathrates.
II Methods
II.1 Clathrate model
To study the vibrational structure and thermal transport properties of type-I clathrates, we use a coarse-grained model of the material [8], illustrated in Fig. 1b. We model dodecahedral and tetrakaidecahedral cages with single “hollow” atoms, which are arranged on an FCC lattice and interact through a Lennard-Jones (LJ) potential. Guest atoms are described by smaller atoms at each FCC site, which interact with the cage atoms at those sites through anharmonic, quartic potentials. The Hamiltonian for our clathrate model is given by
$\begin{split}H&=\frac{1}{2}\sum_{\bm{m}\alpha}\left(\frac{\bm{p}_{\bm{m}\alpha}^{2}}{m_{\alpha}}\right)+\frac{1}{2}\sum_{\bm{m}\alpha,\bm{n}\beta}^{\text{cage}}V_{\text{LJ}}\big{(}\left|\bm{r}_{\bm{m}\alpha}-\bm{r}_{\bm{n}\beta}\big{|}\right)\\ &\hskip 10.00002pt+\sum_{\bm{m}}\sum_{\alpha}^{\text{cage}}\sum_{\beta}^{\text{guest}}{}^{{}^{\prime}}V_{\text{q}}^{\alpha\beta}\left(\bm{r}_{\bm{m}\alpha}-\bm{r}_{\bm{m}\beta}\right)\end{split}$ (2)
$V_{\text{LJ}}\left(r\right)=4\epsilon\left[\left(\frac{\sigma}{r}\right)^{12}-\left(\frac{\sigma}{r}\right)^{6}\right]$ (3)
$V_{\text{q}}^{\alpha\beta}\left(\bm{r}\right)=\sum_{i}\left(\frac{1}{2}K_{\beta,i}r_{i}^{2}+g_{\beta,i}r_{i}^{4}\right)\,,$ (4)
where the primed summation indicates that only cages and guests on the same lattice site interact. Here, $\bm{m},\bm{n}$ are lattice translation vectors, $\alpha,\beta$ are indices over atoms in the unit cell, and $i$ is an index over the Cartesian directions. The position, momentum, and mass of atom $\alpha$ in cell $\bm{m}$ are given by $\bm{r}_{\bm{m}\alpha}$, $\bm{p}_{\bm{m}\alpha}$, and $m_{\alpha}$, respectively. Further details about our model, including parameters for the LJ and quartic potentials, are given in our previous work [8]. Importantly, the symmetry of our model implies that cubic force constants involving guest atoms are identically zero, i.e., $\partial^{3}\mathcal{V}/\partial u_{\bm{m}\alpha i}\partial u_{\bm{n}\beta j}\partial u_{\bm{l}\gamma k}=0$, where $\alpha,\beta$ and/or $\gamma$ are guest atoms, although this is not true of the fully atomic material. This difference has important implications for the performance of perturbation theory and comparison to previous works, as discussed below.
II.2 Anharmonic lattice dynamics with vibrational dynamical mean-field theory
Figure 1: (a) Crystal structure of BaGG [17], where Ba(1) atoms are in dodecahedral cages (gold) and Ba(2) atoms are in tetrakaidecahedral cages (blue). (b) Schematic of coarse-grained model of filled clathrates BaGG and SrGG, where the large cage atoms are colored according to their quartic cage-guest potentials. (c) Spectral functions of BaGG (left) and SrGG (right) at 300 K calculated using SCP+VDMFT. Black dotted lines indicate the harmonic dispersion relation, and black dashed lines indicate the renormalized dispersion calculated using SCP. (d) The spectral function of SrGG at 300 K at $\bm{k}=[\xi 00]$, where $\xi=\pi/a$, calculated using MD, VDMFT, and SCP+VDMFT (top). Anharmonic frequency shifts (bottom left) and linewidths (bottom right) of SrGG at 300 K obtained from the self-energy calculated using MD, SCP+VDMFT, and SCP+PT sampled on a $4\times 4\times 4$ Gamma-centered grid of the BZ. The frequency shift is calculated as $\Delta\omega_{\lambda}(\bm{k})=\omega_{\text{eff},\lambda}(\bm{k})-\omega_{\lambda}(\bm{k})$, where $\omega^{2}_{\text{eff},\lambda}(\bm{k})=\omega^{2}_{\lambda}(\bm{k})+2\omega_{\lambda}(\bm{k})\mathrm{Re}\bm{\pi}_{\lambda,\lambda}\big{(}\bm{k},\omega_{\text{eff},\lambda}(\bm{k})\big{)}$ and was solved for iteratively. The linewidth is calculated as $\Gamma_{\lambda}(\bm{k})=2\omega_{\lambda}(\bm{k})\mathrm{Im}\bm{\pi}_{\lambda,\lambda}\big{(}\bm{k},\omega_{\text{eff},\lambda}(\bm{k})\big{)}/\omega_{\text{eff},\lambda}(\bm{k})$.
In VDMFT, we calculate the anharmonic phonon Green’s function (GF) of the periodic lattice [18],
$D_{\lambda,\lambda^{\prime}}(\bm{k},t)=-i\theta(t)\langle\left[u_{\lambda}(\bm{k},t),u_{\lambda^{\prime}}(-\bm{k},0)\right]\rangle\,,$ (5)
[…]
$u_{\lambda}(\bm{k})=\sum_{\alpha i}c_{\alpha i,\lambda}(\bm{k})u_{\alpha i}(\bm{k})\,,$ (6)
expressed in the basis of of translational symmetry adapted, mass-weighted atomic displacements of atom $\alpha$ in Cartesian direction $i$,
$u_{\alpha i}(\bm{k})=N^{-1/2}\sum_{\bm{m}}e^{-i\bm{k}\cdot\bm{R}_{\bm{m}\alpha}}\sqrt{m_{\alpha}}u_{\bm{m}\alpha i}\,,$ (7)
where $\bm{R}_{\bm{m}\alpha}$ is the equilibrium position of atom $\alpha$ in cell $\bm{m}$. The GF satisfies a Dyson equation,
$\bm{D}^{-1}(\bm{k},\omega)=\bm{D}_{0}^{-1}(\bm{k},\omega)-2\bm{\Omega}(\bm{k})\bm{\pi}(\bm{k},\omega)\,,$ (8)
where $\bm{D}_{0}(\bm{k},\omega)=[\omega^{2}\bm{1}-\bm{\Omega}^{2}(\bm{k})]^{-1}$ is the harmonic GF of the lattice with phonon frequencies $\omega_{\lambda}(\bm{k})$, and $\bm{\pi}(\bm{k},\omega)$ is the self-energy. In our previous work [8], the phonon mode coefficients $c_{\alpha i,\lambda}(\bm{k})$ and frequencies $\omega_{\lambda}(\bm{k})$ were defined as eigenvectors and eigenvalues of the dynamical matrix in the harmonic approximation,
$\bm{\mathcal{D}}_{\alpha i,\beta j}\left(\bm{k}\right)=\frac{1}{\sqrt{m_{\alpha}m_{\beta}}}\sum_{\bm{m}}e^{i\bm{k}\cdot\left(\bm{R}_{\bm{m}\alpha}-\bm{R}_{\bm{0}\beta}\right)}\frac{\partial^{2}\mathcal{V}}{\partial u_{\bm{m}\alpha i}\partial u_{\bm{0}\beta j}}\,,$ (9)
where derivatives of the lattice potential, $\mathcal{V}$, with respect to atomic displacements, $u_{\bm{m}\alpha i}=r_{\bm{m}\alpha i}-R_{\bm{m}\alpha i}$, are evaluated at the equilibrium lattice configuration. Here, we improve this harmonic description and use self-consistent phonon theory (SCP) [19, 20, 21, 22, 23] to obtain a more accurate temperature-dependent quasiparticle description of phonons. In our SCP calculations, the phonon modes and frequencies are defined via eigenvectors of a modified dynamical matrix, $\bm{\mathcal{D}}(\bm{k})+\bm{\mathcal{W}}(\bm{k})$, where $\bm{\mathcal{W}}(\bm{k})$ is the mean-field contribution due to quartic anharmonicity.
In VDMFT, the dynamics of the periodic lattice are mapped onto an “impurity problem” that consists of a single unit cell coupled to a fictitious bath of harmonic oscillators through a tailored spectral density [24, 25, 26, 27, 15, 8]. The local self-energy in a single unit cell, $\bm{\pi}(\omega)$, is calculated through the exact solution of the impurity problem and is used to approximate the self-energy of the lattice GF, $\pi_{\alpha i,\beta j}(\bm{k},\omega)\approx\pi_{\alpha i,\beta j}(\omega)$, thus providing a nonperturbative treatment of anharmonicity. The lattice GF and impurity problem are iteratively updated until self-consistency is achieved. We perform SCP calculations using the ALAMODE package [28] (details given in Supplemental Material, SM) and use that quasiparticle basis in combination with VDMFT, following the approach described in detail in our previous work [15, 8]. As in our previous work, the impurity GF is calculated by neglecting nuclear quantum effects and treating the nuclei classically.
Figure 1c illustrates the trace of the anharmonic spectral functions, $\bm{A}(\bm{k},\omega)=-\pi^{-1}\mathrm{Im}\bm{D}(\bm{k},\omega)$, of BaGG and SrGG calculated using SCP+VDMFT at 300 K. Anharmonicity causes hardening of all phonon modes, especially those flat modes dominated by guest atoms rattling in the tetrakaidecahedral cages, which increase in frequency by 3 cm-1 and 12 cm-1 for BaGG and SrGG, respectively [8]. These frequency shifts are accurately captured by the SCP quasiparticle basis, as shown by the black dashed lines in Fig. 1c. Finite lifetimes due to phonon-phonon scattering cause broadening of the quasiparticle dispersion, which is captured by the SCP+VDMFT spectral function. Again, the guest-dominated modes display significant broadening; in particular, the Sr rattling mode acquires a linewidth of 8 cm-1, which corresponds to a short lifetime of 0.67 ps. The hardening and broadening of the rattling modes in these filled clathrates also affects their avoided crossing with the cage acoustic modes, thereby impacting the thermal conductivity of these materials.
We assess the accuracy of our SCP+VDMFT approach by comparing the spectral function of SrGG, the most anharmonic system studied here, with the exact spectral function computed using molecular dynamics (MD) simulations of a large supercell with periodic boundary conditions. The spectral functions at the $X$ point of the BZ are illustrated in Fig. 1d along with the frequency shifts and linewidths calculated on a 4$\times$4$\times$4 grid of the BZ, demonstrating excellent agreement between SCP+VDMFT and MD. We also compare linewidths to those calculated with lowest-order perturbation theory (PT) of three-phonon scattering processes using the SCP quasiparticle basis, which is the conventional approach for computing phonon linewidths [5]. Details regarding the MD and SCP+PT calculations are given in the SM. Unlike MD and SCP+VDMFT, Fig. 1d shows that SCP+PT inaccurately predicts negligible broadening for all modes in SrGG. This failure of SCP+PT has significant implications for calculations of the thermal conductivity, which rely on the accurate description of phonon linewidths, as described below.
II.3 Thermal conductivities
II.3.1 Green-Kubo formalism
The thermal conductivity, $\bm{\kappa}$, relates the macroscopic heat flux in a system, $\bm{J}$, with an applied temperature gradient. In the linear response regime, the thermal conductivity can be computed using the Kubo formula [29], which depends on the autocorrelation function of the heat-flux operator:
$\kappa^{ij}$ $=\frac{V}{T}\lim_{\omega\rightarrow 0}\frac{\mathrm{Im}\chi^{ij}(\omega+i\eta)}{\omega}$ (10a)
$\chi^{ij}(t)$ $=\frac{i}{\hbar}\theta(t)\langle\left[J^{i}(t),J^{j}(0)\right]\rangle$ (10b)
where $V$ is the supercell volume. In the classical limit, this reduces to the familiar expression
[…]
but here we aim to maintain quantum statistics, which is important for the low temperature conductivity. In atomistic or lattice models, the heat flux operator is not uniquely defined. Here, we use the Hardy definition of the harmonic heat-flux operator [30],
$\begin{split}J^{i}&=\frac{\hbar}{2V}\sum_{\bm{k}}\sum_{\lambda,\lambda^{\prime}}\omega_{\lambda}(\bm{k})v^{i}_{\lambda,\lambda^{\prime}}(\bm{k})\\ &\hskip 10.00002pt\times\left[a^{\dagger}_{\lambda}(\bm{k})+a_{\lambda}(-\bm{k})\right]\left[a_{\lambda^{\prime}}(\bm{k})-a^{\dagger}_{\lambda^{\prime}}(-\bm{k})\right]\,,\end{split}$ (12)
where $a^{\dagger}_{\lambda}(\bm{k})$ and $a_{\lambda}(\bm{k})$ are the creation and annihilation operators, respectively, of mode $\lambda$, and $\omega_{\lambda}(\bm{k})$ is its frequency. The generalized velocity matrix [16] is
$\bm{v}_{\lambda,\lambda^{\prime}}(\bm{k})=\frac{1}{2\sqrt{\omega_{\lambda}(\bm{k})\omega_{\lambda^{\prime}}(\bm{k})}}$
×$\quad\times\Bigg{(}\sum_{\alpha i,\beta j}c_{\alpha i,\lambda}(\bm{k})\nabla_{\bm{k}}[\bm{\mathcal{D}}(\bm{k})+\bm{\mathcal{W}}(\bm{k})]_{\alpha i,\beta j}c^{*}_{\beta j,\lambda^{\prime}}(\bm{k})\Bigg{)}\,,$ (13)
where, in the absence of degeneracies, the diagonal elements coincide with the usual phonon group velocities, $\bm{v}_{\lambda,\lambda}=\nabla_{\bm{k}}\omega_{\lambda}(\bm{k})$.
While the response function, $\chi_{ij}(t)$, depends on products of four bosonic operators, or a two-particle GF, it can be approximated as a product of two single-particle GFs by neglecting vertex corrections [31]. This approximation, which is exact in infinite dimensions [24], is expected to be valid for materials with low thermal conductivity, in which resistive, Umklapp scattering processes are dominant over momentum-conserving, normal scattering processes that are characteristic of the hydrodynamic regime of thermal transport [32]. This approximation yields the thermal conductivity [16],
$\kappa^{ij}=\frac{\pi\hbar}{VT}\sum_{\bm{k}}\sum_{\lambda,\lambda^{\prime}}\omega_{\lambda}(\bm{k})\omega_{\lambda^{\prime}}(\bm{k})v^{i}_{\lambda,\lambda^{\prime}}(\bm{k})v^{j}_{\lambda^{\prime},\lambda}(\bm{k})$
×$\qquad\times\int_{-\infty}^{\infty}d\omega\left[\omega^{2}+\omega^{2}_{\lambda}(\bm{k})\right]A_{\lambda,\lambda}(\bm{k},\omega)A_{\lambda^{\prime},\lambda^{\prime}}(\bm{k},\omega)$
×$\qquad\qquad\qquad\times\Bigg{[}-\frac{\partial n_{T}(\omega)}{\partial\omega}\Bigg{]}\,,$ (14)
where $n_{T}(\omega)=\big{(}e^{\hbar\omega/k_{B}T}-1\big{)}^{-1}$ is the Bose-Einstein distribution. Equation (14) contains both intraband ($\lambda=\lambda^{\prime}$) and interband ($\lambda\neq\lambda^{\prime}$) contributions [33, 34, 16]. The interband contributions become important in materials for which phonon broadening is on the same order as energy spacing between phonon bands, as seen in the clathrates studied here, especially SrGG (Fig. 1c).
II.3.2 Boltzmann transport equation within the single mode approximation
Figure 2: Intraband contributions to the thermal conductivities of GG (left), BaGG (center), and SrGG (right) calculated at 300 K using the Green-Kubo formalism with spectral functions calculated using SCP+VDMFT. Points indicate the mode-resolved intraband thermal conductivity, where the size and color of each point correspond to the relative magnitude. Black lines indicate the SCP dispersion at 300 K. The top left of each panel gives the value of the total intraband thermal conductivity at 300 K.
The full expression for the thermal conductivity given by Eq. (14) can be simplified under the assumption that the phonon quasiparticle picture is valid such that interband transitions are negligible and that the phonon spectral densities have Lorenztian lineshapes centered at the harmonic phonon frequencies $\omega_{\lambda}(\bm{k})$ with widths $\Gamma_{\lambda}(\bm{k})$. These approximations yield an expression for the thermal conductivity that is identical to the BTE-SMA expression in Eq. (1), with $v_{\lambda}^{i}(\bm{k})=v_{\lambda,\lambda}^{i}(\bm{k})$ and $\tau_{\lambda}(\bm{k})=1/2\Gamma_{\lambda}(\bm{k})$.
The reliability of this BTE-SMA approach depends on the accuracy of the above approximations and of the frequencies and lifetimes used. For example, the frequencies and lifetimes can be calculated by a combination of mean-field and perturbation theory, as is commonly done, or extracted from our SCP+VDMFT approach, as shown in Fig. 1d. Results obtained using both approaches will be compared in the following sections.
III Results and Discussion
III.1 Intraband transport
First, we examine the intraband contributions to the thermal conductivity. We calculate the thermal conductivity at 300 K using the Green-Kubo formalism [Eq. (14)] with anharmonic spectral functions calculated with SCP+VDMFT sampled on a 12$\times$12$\times$12 Gamma-centered grid of the BZ and consider only the terms where $\lambda=\lambda^{\prime}$. The intraband thermal conductivity of the empty clathrate, GG, is 35.2 W/mK, which is more than 20 times greater than that of the filled clathrates BaGG and SrGG, which have values of 1.24 W/mK and 0.70 W/mK, respectively. BaGG features weaker cage-guest interactions than SrGG, indicating the role of cage-guest anharmonicity in reducing the thermal conductivity.
Figure 2 illustrates the SCP dispersions of GG, BaGG, and SrGG at 300 K along a high-symmetry path through the BZ as well as each phonon mode’s contribution to the intraband thermal conductivity. In GG, acoustic modes have the largest contributions to the thermal conductivity due to both their high phonon mode velocities and relatively long lifetimes. Optical modes contribute, as well, especially those near the zone edges with larger group velocities. In the filled clathrates, flat optical modes corresponding to rattling motions of X(2) guest atoms inside 24-atom cages, where X=Ba,Sr, cut through the acoustic modes of the cage lattice, leading to the avoided crossings that are characteristic of these materials [8]. Consequently, the acoustic modes of the filled clathrates have diminished contributions to the thermal conductivity, especially as they approach the avoided crossing near the BZ edges. As seen in Fig. 2, points corresponding to the flat, guest-dominant modes have negligible contributions to the thermal conductivity.
The SM shows the calculated mean-free-paths of acoustic modes of GG, BaGG, and SrGG at various temperatures. At all temperatures, the filled clathrates have mean-free-paths that are an order of magnitude or more smaller than those of the empty clathrate. This reduction is due to a sharp decrease in phonon lifetimes for the filled clathrate acoustic modes in a wide frequency range, confirming the role of nonresonant scattering in these materials [4]. In contrast, the acoustic phonon mode velocities are unchanged near the BZ center, and they only decrease as they approach the zone edge and gain significant character of the X(2) guest atoms.
III.2 Interband transport
Figure 3: (a) Interband contributions to the thermal conductivities of GG, BaGG, and SrGG calculated at different temeratures using the Green-Kubo formalism. (b) Interband thermal conductivities of SrGG showing transport between modes with frequencies $\omega_{\lambda}$ and $\omega_{\lambda\prime}$.
Due to the strong anharmonicity and significant mode mixing of the filled clathrates, especially near the avoided crossing of the cage-acoustic and guest-optical modes, we investigate the role of interband transport in the overall thermal conductivity [33, 35]. Figure 3a shows the percent of the total thermal conductivity that is due to interband transport. As expected, GG has negligible interband transport at all temperatures because of its relatively harmonic lattice dynamics, for which the particle-like picture of thermal transport is dominant. The filled clathrates, however, have significant contributions from interband transport that increases with increasing temperature. At 600 K, 8% and 11% of the total thermal conductivity for BaGG and SrGG, respectively, are due to interband transport.
To better understand the mechanism underlying interband transport in SrGG, we analyze the mode-resolved interband contributions to the thermal conductivity at different temperatures, which are illustrated in Fig. 3b. There are few interband transitions at 50 K, and they increase slightly at 100 K, primarily occurring between phonon modes that are close in energy (i.e., $\omega_{\lambda}\sim\omega_{\lambda\prime}$). At 300 K, interband transport begins to occur more significantly between modes that are further apart in energy. In particular, a plus-shaped pattern appears around $\omega_{\lambda}=\omega_{\lambda\prime}\sim 32\,$cm-1, which is the frequency of the flat rattling Sr(2) modes at 300K. These modes have broad linewidths of 8 cm-1 and and are intersected by several of the cage-acoustic modes (Fig. 1c). The overlapping spectral densities of the modes along these regions enable significant interband transport. Another plus-shaped region appears around 60 cm-1, which corresponds to the overlapping of the cage-optical and higher-energy X(2)-dominant modes at that energy. Further interband transport between modes with larger energy differences occurs at 600 K, and additional plus-shaped regions appear around the X(1) guest modes at 80 cm-1. As the X(2) rattling mode continues to harden and broaden at increased temperatures, the interband transitions around that frequency range become more prominent.
This analysis directly demonstrates the role of guest atoms in enabling interband transport, which is negligible in the empty clathrates. However, we emphasize that while a significant percentage of the total thermal conductivity is due to interband transport enabled by guest-dominant modes, the absolute value of the interband thermal conductivity is small (the interband thermal conductivity of BaGG and SrGG at 300 K is 0.051 W/mK and 0.053 W/mK, respectively), and the primary consequence of the guest atoms are to reduce the overall thermal conductivity. Despite the differences in anharmonicity between BaGG and SrGG, which is reflected by differences in their spectral functions (Fig. 1c), they have similar values for interband transport at 300 K. While the broad SrGG rattling modes have more spectral overlap with acoustic modes than those of BaGG, which enables interband transfer, they also have shorter lifetimes and lower frequencies, which reduce those transitions’ contributions to the thermal conductivity. These results indicate that a complex balance of factors and spectral features underlie the thermal conductivity, which is characterized by a single value; although brute force molecular dynamics is commonly applied to calculate the thermal conductivity as the integral of the heat flux autocorrelation function [Eq. (11)], such an approach lacks the atomistic insights available from the mode-resolved Green’s function approach taken here.
III.3 The role of nonperturbative effects on thermal conductivity
Figure 4: (a) Thermal conductivities of GG (blue), BaGG (green), and SrGG (red) calculated using the Green-Kubo formalism [Eq. (14)] with spectral functions calculated using SCP+VDMFT (triangles) and using the BTE formalism [Eq. (1)] with lifetimes calculated using SCP+PT (circles). All calculations are sampled on a $12\times 12\times 12$ Gamma-centered grid of the BZ. (b) Thermal conductivities at higher temperatures, where black points are experimental measurements from Ref. [11]. Grey lines indicating different power laws are a guide to the eye.
To better understand the role of nonperturbative effects on the thermal conductivity, we compare the thermal conductivity across materials and temperatures computed using two different methods: the Green-Kubo formalism [Eq. (14)] with spectral functions calculated using SCP+VDMFT (GK-VDMFT) and the BTE-SMA approach [Eq. (1)] with SCP phonons and lifetimes calculated using conventional lowest-order PT of the three-phonon scattering processes (BTE-PT). Details regarding the BTE-PT calculations are provided in the SM.
When calculated using GK-VDMFT, the thermal conductivity for all materials peaks around 25-50 K and then decreases with increasing temperature. As discussed above, the introduction of guest atoms decreases the thermal conductivity by more than one order of magnitude, and BaGG, which features weaker cage-guest anharmonicity, has a larger thermal conductivity than SrGG at all temperatures. The BTE-PT approach results in thermal conductivity values that are consistent with GK-VDMFT for the quasi-harmonic GG, but they disagree significantly for the filled clathrates. BTE-PT predicts only a factor of two reduction between GG and BaGG, and it incorrectly predicts that the more anharmonic SrGG has a larger thermal conductivity than BaGG. These discrepancies are a result of the failure of SCP+PT to capture the short lifetimes of the filled clathrates (Fig. 1d), causing a significant overestimation of their thermal conductivities. As shown in the SM, using the BTE-SMA formalism with lifetimes determined from SCP+VDMFT calculations yields values that closely align with those calculated by the GK-VDMFT approach, suggesting that the success of the BTE-SMA depends strongly on the quality of the phonon lifetimes used and, in this case, is not due to the exclusion of interband transport. The SM also illustrates comparisons to thermal conductivities calculated according to the classical analog of Eq. (10) using molecular dynamics simulations, showing consistency with the GK-VDMFT results. These comparisons suggest that the ultralow thermal conductivities of the filled clathrates result from four-phonon and higher-order scattering processes that are not described by SCP+PT but that are inherently described by our SCP+VDMFT approach, highlighting the importance of nonperturbative anharmonic effects on observables like the thermal conductivity. In Fig. 4b, we compare our calculated values of the thermal conductivity for BaGG to experimental data, focusing on high-temperature data, where intrinsic effects of anharmonicity dominate over effects due to grain boundary scattering, dynamical disorder, and nuclear quantum effects. We see excellent agreement, validating our approach. At lower temperatures, phonon-grain boundary scattering can become important [4], as discussed in the SM.
Finally, we consider the temperature dependence of the total thermal conductivity (Fig. 4b), as strong interactions can produce anomalous behavior, as in electronic transport properties [36]. Here, we focus on the high-temperature dependence, as our calculations neglect nuclear quantum effects and grain boundary scattering, which may affect the behavior of thermal conductivities at low temperatures. For all materials, BTE-PT predicts a $T^{-1}$ temperature dependence of the thermal conductivity. This temperature dependence is dictated by that of the phonon lifetimes, for which a $T^{-1}$ scaling is inherent to the PT formalism at temperatures greater than the phonon frequency (see SM). In contrast, the GK-VDMFT approach predicts different high-temperature behavior for the different clathrates. BaGG has an intermediate scaling of $\sim T^{-0.9}$, and SrGG, the most anharmonic system, has the most moderate scaling, $\sim T^{-0.6}$.
Figure 5: (a) Renormalized dispersion of acoustic modes near the BZ center of BaGG (left) and SrGG (right) calculated using SCP at 300 K. (b) The acoustic mode lifetimes of BaGG (left) and SrGG (right) extracted from the self-energy calculated using the SCP-VDMFT approach. Grey lines indicating a $T^{-1}$ power law are a guide to the eye.
The high-temperature dependence of the thermal conductivity primarily arises from that of the acoustic phonon lifetimes. Thus, we analyze the temperature dependence of longitudinal and transverse acoustic phonon modes near the BZ center of BaGG and SrGG, as illustrated in Fig. 5. While the mode lifetimes of BaGG largely follow a $T^{-1}$ behavior, the mode lifetimes of SrGG deviate significantly from the $T^{-1}$ behavior that is predicted by perturbation theory and begin to saturate to a milder scaling at 50 K. This high-temperature saturation of the phonon lifetimes, and thus thermal conductivities, is reminiscent of “resistivity saturation” in metals [36, 37, 38, 39] and is a hallmark of nonperturbative interactions. This result indicates that the acoustic modes of the SrGG cage framework are subject to multiphonon scattering processes that are not captured by conventional perturbation theory. Despite their relatively narrow linewidths and long lifetimes (Fig. 1), these acoustic phonons’ low energies and strong coupling to Sr(2) rattling modes makes them more likely to participate in multiphonon scattering processes. Thus, we find that including nonperturbative descriptions of anharmonicity is essential in accurately capturing thermal transport properties, especially at higher temperatures where multiphonon processes are activated.
IV Conclusions
In conclusion, we have developed a framework based on VDMFT for the efficient and accurate calculation of thermal transport properties in real materials, including nonperturbative anharmonic effects and both intraband and interband transport mechanisms. We have applied this methodology to understand the temperature-dependent thermal conductivities of type-I clathrate solids. Although we have used a coarse-grained representation of the atomic structure of these type-I clathrates, our results for the thermal conductivity are qualitatively and quantitatively consistent with ab initio calculations. The authors of Ref. 5 calculate a value of 0.97 W/mK for BaGG at 300 K using the BTE-SMA approach with lifetimes extracted from SCP+PT calculations and ignoring interband transport, and the authors of Ref. 35 calculate values of 1.26 W/mK and 0.77 W/mK for BaGG and SrGG, respectively, using a similar approach while also including interband transport. Those works successfully capture the reduction in thermal conductivities of the filled clathrates, even while using conventional PT, because of the atomistic detail in their clathrate structures. In the atomistic configuration (Fig. 1a), the spherical asymmetry of the cage structures enable three-phonon processes that are strictly zero in our spherically symmetric coarse-grained structure (Fig. 1b). However, results with our coarse-grained model do confirm that the mechanism for thermal conductivity reduction in the filled clathrates is due to nonresonant scattering that decreases lifetimes of cage acoustic modes [4, 5, 35], even at frequencies away from those of the guest rattling modes, and that changes in acoustic mode velocity play a relatively minor role.
Finally, as shown in Fig. 4, our calculations are able to capture the experimental high-temperature behavior of BaGG, while the results of Refs. 4, 5 show a steeper scaling with increasing temperature. This discrepancy may be due to role of nonperturbative effects discussed above.
The role of nonperturbative effects has also been suggested by Ikeda et al.[6] as being important for capturing the experimental high-temperature dependence of thermal conductivity. However, Ikeda et al. also discuss that the thermal conductivity peaks at 10 K, a temperature lower than that expected by the traditional Debye model. We calculate thermal conductivities that peak between 25-50 K, but we neglect nuclear quantum effects, which may be important at these low temperatures.
Future work includes applying our VDMFT framework to a fully atomistic description of the type-I clathrate materials to evaluate the relative importance of three-phonon versus multi-phonon scattering processes as well as including quantum nuclear dynamics to understand their effect on thermal conductivities, especially at very low temperatures.
Acknowledgements. This work was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, under Award No. DESC0023002. We acknowledge computing resources from Columbia University’s Shared Research Computing Facility project, which is supported by NIH Research Facility Improvement Grant No. 1G20RR030893-01, and associated funds from the New York State Empire State Development, Division of Science Technology and Innovation (NYSTAR), Contract No. C090171.